Find the general solution of the first-order linear partial differential equation

xux yuy = u and investigate first-order, quasi-linear and linear partial differential equations


1
Expert's answer
2021-02-24T07:24:06-0500

We have the linear differential equation


"xu_x+yu_y=u"


The system of eqautions is,

"\dfrac{dx}{x}=\dfrac{dy}{y}=\dfrac{du}{u}"


Now, taking first two terms

"\dfrac{dx}{x}=\dfrac{dy}{y}"


Integrating Both the sides-


"lnx=lny+C1\\

lnx-lny=\text{ln}e^{C1}"


"(\dfrac{x}{y})=e^{C1}\\

\dfrac{x}{y}=C1'\\

\phi=\dfrac{x}{y}=C1'"


Now, taking last two terms

"\dfrac{dy}{y}=\dfrac{du}{u}\\

lny=lnu+C2\\

lny=lnu+lne^{C2}\\\\

ln(\dfrac{y}{u})=lne^{C2}"


"\dfrac{y}{u}=e^{C2}"


"\dfrac{y}{u}=C2'"


"\epsilon=\dfrac{y}{u}=C2'"


Hence, the general solution is "f(\phi,\epsilon)=f(\dfrac{x}{y},\dfrac{y}{u})=0"